Temperature of a wire is increased by 3°C, when a current is passed in the wire for sometime. If the current is doubled , then by what amount will the temperature of wire increase during same interval of time.

Grade:10

1 Answers

Arun
6 years ago
Dear Sneha
As we know
H = I² Rt
Hence
H1/H2 = (I1)²/(I2
H2/H1 = 4
Now
Theta proportional to H
Now
θ21 = H2/H1
θ2 = 4 * 3 = 12 °c
Regards
Arun (askIITians forum expert)